OpenStudy (anonymous):

d/dx A(1+x^2)^1/2 where A is constant

OpenStudy (stamp):

\[d(a\sqrt{1+x^2})/dx=\frac{ax}{\sqrt{1+x^2}}\]Cch-chchainrule

OpenStudy (anonymous):

ohh, so, i still need to use chain rule for this ? because i thought the constant can be consider as a number

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

-_- now only i see this, hahaha, thanks for helping :D